Two capacitors give an equivalent capacitance of 9.00 x10^(-6)F when connected in parallel and an equivalent capacitance of 2.00 x10^(-9) F when connected in series. What is the capacitance of each capacitor?
1
Expert's answer
2021-02-18T18:56:43-0500

The equivalent capacitances of two capacitors connected in series and in parallel can be written as follows:


"C_{eq,s}=\dfrac{C_1C_2}{C_1+C_2},""C_{eq,p}=C_1+C_2."

Let's express "C_1" from the last equation in terms of "C_{eq,p}" and "C_2":


"C_1=C_{eq,p}-C_2."

Substituting "C_1"into the first equation and simplifying, we get:


"C_2^2-C_{eq,p}C_2+C_{eq,s}C_{eq,p}=0,"

Let's substitute the numbers and solve the quadratic equation:


"C_2^2-9.0\cdot10^{-6}C_2+18\cdot10^{-15}=0."

This quadratic equation has two roots: "C_1=9\ \mu F" and "C_2=2\ nF". Therefore, the capacitance of each capacitor: "C_1=9\ \mu F, C_2=2\ nF."

Answer:

"C_1=9\ \mu F, C_2=2\ nF."